OpenStudy (anonymous):

Suppose that f(x)=sqrt(x+4), g(x)=x^2+1, and h(x)= x+2 Find (f∘g∘h)(x) =

OpenStudy (amistre64):

f(g(h(x))) = f(h(x)^2+2) = sqrt((h(x)^2+2)+4)=sqrt((x+2)^2+2+4)

OpenStudy (amistre64):

almost got that, typoed the 2

OpenStudy (amistre64):

f(g(h(x))) = f(h(x)^2+1) = sqrt((h(x)^2+1)+4)=sqrt((x+2)^2+1+4) is more like it

OpenStudy (anonymous):

thanks!

OpenStudy (amistre64):

yep, those compund stuff can be a headache if you dont keep your wits about ya